The length of a rectangle is 5 more than twice the width. The perimeter is 130. What is the area?
Answer by checkley71(8403)

L=2W+5
PERIMETER=2L+2W
130=2(2W+5)+2W
130=4W+10+2W
130=6W+10
6W=130-10
6W=120
W=120/6
W=20 ANSWER FOR THE WIDTH.
L=2*20+5
L=40+5
L=45
PROOF
2*45+2*20=130
90+40=130
130=130